1) The interest on $120 for 8months is $3.60, find the rate % p.a

2) Pie = 3.14159, e= 2.71828, find pie^2 ÷ e^3 to 3 figures

3) Solve (2x-1)(8x+3)= (4x-1)^2


1
Expert's answer
2021-12-07T09:38:28-0500

a)


"I=P(1+\dfrac{r}{12})^8-P"

"3.60=120((1+\dfrac{r}{12})^8-1"

"(1+\dfrac{r}{12})^8=1.03"

"r=12((1.03)^{1\/8}-1)"

"r\approx0.04442"

"4.442\%"


2)


"\dfrac{\pi e^2}{e^3}=\dfrac{\pi}{e}\approx \dfrac{3.14159}{2.71828}\approx1.16"

3)


"(2x-1)(8x+3)=(4x-1)^2"

"16x^2+6x-8x-3=16x^2-8x+1"

"6x=4"

"x=\dfrac{2}{3}"